OpenStudy (anonymous):

Let f(x)= x^2-1 and g(x)=2x-3 Find (g * f) (5)

OpenStudy (lgbasallote):

(g*f)(5) = g(5) * f(5) g(5) = 2(5) - 3 = 10 - 3 = 7 f(5) = (5)^2 - 1 = 25 - 1 = 24 so g(5) * f(5) = 7 x 24 got it?

OpenStudy (anonymous):

Thats the answer i got also but there is not answer choice 168. I have -45 ,48, 5 and 45 as the choices. help ??!!

OpenStudy (lgbasallote):

wait...did it look like something like \[(f \quad o \quad g)(5)\] or \[(f \cdot g)(5)\]

OpenStudy (anonymous):

(f o g ) (5)

OpenStudy (anonymous):

If I recall that's the same as f ( g (x) ), nested functions basically (one inside the other). Formally it's pronounced: "f composed with g of x" or "f of g of x". f(x)= x^2-1 and g(x)=2x-3 You input g(x) into f(x), and then put 5 in for x Step 1: f(g(x))= (2x-3)^2-1 Step 2: f(g(5))= (2(5)-3)^2-1

OpenStudy (anonymous):

See if you get the answer now, and watch the order of operations here :-)

OpenStudy (lgbasallote):

sorry i had technnical difficulties but @agentx5 just said how to do it..

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

Im so lost. lol. Thanks for trying to help me.

OpenStudy (anonymous):

Yes. thats the answer choices.